Operations Report – April 24th

Crew 211 Operations Report 24-APR-2019

SOL: 3

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10:30

On (last night): 8:30 pm

Off (this morning): 7 am

Solar— SOC 100 % at 20:10 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~240 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 110 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6 units

Toilet tank emptied: yes

Deimos rover used: no (rover is out of order according to Director)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 89.0h

Beginning charge: 100

Ending charge: 84.4

Currently charging: yes

Opportunity rover used: yes

Hours: 57.6

Beginning charge: 95

Ending charge: 74

Currently charging: no

Curiosity rover used: yes

Hours: 90.8h

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Notes on rovers: Yesterday during the EVA the rover Opportunity suffered overheating and could not be started back. This morning’s mission was to bring it back so that the mechanics could fix it.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some basil and tomatoes were harvested for cooking.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Simon finished his UWB stands and Carl continued assembling more probes during the afternoon after children finished visiting us.

Summary of any observatory issues: The Astronomer’s computer encounters sometimes freeze and displays artifacts on the screen. This might be the beginning of a GPU failure.

Summary of health and safety issues: The member with the twisted ankle feels better and can walk, but for extra safety HSO does not recommend EVA.

Questions, concerns and requests to Mission Support:

* We had an Internet problem last night which prevented us to send the operations report to CapComm. It solved itself at around 10am this morning.

* We overestimated the static water tank capacity, and after checking the level properly with a rope and measuring volumes we reestimated its current level. We will try to diminish consumption as much as we can, but it is already very low and we are investigating other reasons.

* The water pump leak is at a plastic joint. To fix it we would need to stop the waterflow and replace the joint with a new one. We will try to document further the following days. Attaching a picture of the pump with the mail (faulty joint is in red, and you can see that water has been leaking for it).

Copyright © The Mars Society. All rights reserved. | Main Site